Exercise 5-8 Compute the Margin of Safety [LO5-7]

Molander Corporation is a distributor of a sun umbrella used at resort hotels. Data concerning the next month’s budget appear below:

Selling price per unit $ 30
Variable expense per unit $ 17
Fixed expense per month $ 10,530
Unit sales per month 960

Required:

1. What is the company’s margin of safety? (Do not round intermediate calculations.)

2. What is the company’s margin of safety as a percentage of its sales? (Round your percentage answer to 2 decimal places (i.e. .1234 should be entered as 12.34).)

1. Margin of safety (in dollars)
2. Margin of safety percentage %
Answer #1

Contribution margin=Sales-Variable cost

=(30-17)=$13 per unit

Breakeven sales=Fixed cost/Contribution margin

=(10530/13)=810 units

Margin of safety=Total sales-Breakeven sales

=(960-810)=150 units

=(150*30)=$4500

Margin of safety %=Margin of safety/Total sales

=(150/960)=15.63%(Approx).